Solve the system of equations. If the system is inconsistent or has dependent equations, say so.
x + 3y + 2z = 11
4y + 9z = -12
x + 7y + 11z = -11

Answer by josmiceli(6796)

(1) x + 3y + 2z = 11
(2)4y + 9z = -12
(3)x + 7y + 11z = -11
looks inconsistent
subtract (1) from (3)
4y + 9z = -22
but (2) says
4y + 9z = -12
both can't be true